# BrainOmni: A Brain Foundation Model for Unified EEG and MEG Signals

### v1

@@ 4.4 Effectiveness of Sensor Encoder
An ablation study is conducted to assess the effectiveness of the proposed Sensor Encoder in modelling the spatial characteristics of sensors. Apart from directly removing the sensor embedding, we also develop a pure temporal version of BrainTokenizer where multi-channel EMEG data is treated as uncorrelated single-channel data. Results shown in Table 6, it can be observed that the standard BrainOmni consistently outperforms the pure temporal model. And the exclusion of sensor embedding significantly undermines the downstream performance, which demonstrates the effectiveness of the proposed Sensor Encoder.
5 Analysis
@@ 5.1 Pretraining Stability and Codebook Contribution
To analyse the stability when incorporating diverse devices and signals into pretraining, we plot the training curves in Fig. 3(a). Overall, BrainTokenizer converges relatively quickly, and the validation loss remains consistent with the training loss. This demonstrates that the proposed BrainTokenizer training framework generalises well across different devices and signal types. Fig. 3(b) plots the mask prediction performance for each RVQ layer. It can be seen that the first codebook aggregates richer semantic information and achieves higher accuracy in mask prediction, while the prediction accuracy decreases progressively for the subsequent lower-level codebooks.
5.2 Number of Latent Source Variables
To investigate the effect of the number of latent source activities, we trained variants of BrainTokenizer with different numbers of source activities. Results are shown in Fig. 3(c). As the number of source activities increases, the EMEG signal reconstruction loss gradually decreases. When the number of source variables reaches beyond 16, the trend of loss reduction weakened with further increases in the number of latent source activities. To balance the amount of information retained and computational efficiency, we chose 16 as the number of latent source variables.
@@ 5.3 Visualisation of Reconstruction Results
Fig. 4 illustrates the reconstruction performance of BrainTokenizer. We compare the waveforms and topographic maps before and after reconstruction to highlight the model’s ability to capture the spatiotemporal electromagnetic fields of EEG and MEG in both temporal and spatial domains. The results show that the model effectively preserves the main trends and finer details of the waveforms, while smoothing out high-frequency noise spikes. In terms of spatial representation, the reconstructed topographic maps maintain the original activation patterns and structural integrity. Notably, the model also achieves strong reconstruction performance on previously unseen devices and samples, demonstrating robust generalisation across different recording setups.
@@ 6 Conclusion
In this paper, we introduce BrainOmni, the first brain foundation model that generalises across heterogeneous EEG and MEG signals. To model signals from different devices and modalities, we propose BrainTokenizer, which infers spatiotemporal patterns of brain activity from the observed EMEG signals and generates quantised discrete tokens. To address the inherent heterogeneity arising from various recording devices, we develop a flexible Sensor Encoder that leverages physical sensor properties. Through large-scale self-supervised joint pretraining on EEG and MEG data, BrainOmni significantly exceeds the performance of existing foundation models and state-of-the-art task-specific baselines across various downstream tasks, and demonstrates effective generalisation capability to unseen devices and modalities. Furthermore, extensive ablation analysis highlights the consistent benefits of joint EMEG pretraining, underscoring the advantage of unified modelling strategies. We believe BrainOmni represents an important step towards building versatile and scalable foundation models for neural recordings, opening new pathways for unified brain signal representation learning.
@@ Appendix A Broader Impacts and Limitations
As the first foundation model enabling unified analysis across EEG and MEG modalities, BrainOmni represents a crucial step toward generalizable and scalable neural representation learning. BrainOmni, as the first foundation-model support for MEG and joint EMEG analysis, offers substantial application potential in neuroscience research, health monitoring, and clinical diagnosis. Additionally, the model’s enhanced generalisation on heterogeneous devices could reduce the performance degradation typically associated with device and dataset variability.
@@ B.1 Source Current Estimation
[…]
@@
[…]
@@
The inverse problem—estimating the sources $\mathbf{j}(t)$ from the measurements $\mathbf{y}(t)$—is ill-posed and requires regularization. Minimum-norm estimation solves this by seeking the source configuration with the smallest $\ell_{2}$-norm that still explains the data:
@@ B.2 Model Explanation From the Perspective of Neuroscience
Inspired by this conceptual framework, our proposed BrainTokenizer closely mirrors these two steps. Specifically, the BrainTokenizer (excluding the RVQ quantization step) corresponds directly to the backward solution stage, where it maps the raw EMEG signals and sensor device layout into latent source variables through a parametric module. Recall the classical backward solution,
@@
which relies on a fixed linear inverse operator $\mathbf{W}$ constructed from a leadfield metrix $\mathbf{L}$ and a noise covariance matrix $\mathbf{C}_{n}$. While in our architecture, we use the attention weights in cross-attention mechanisms to construct a time-varying parametric inverse operator $\mathbf{W}_{\mathbf{\theta}}(\mathbf{Z}_{\text{time}},\mathbf{L},\mathbf{S})$, where the operator is related to sensor prosperity and the measured data since the different fields recorded outside may indicate that there is a change in the pattern of neuro activity, formulated as:
@@
where $\mathbf{Q}$ is a set of learnable query embeddings that dynamically aggregate temporal and spatial information, $\mathbf{K}$ is the combination of temporal representation $\mathbf{Z}_{\text{time}}$ from SEANet encoder and sensor embeddings $\mathbf{V}(\mathbf{L},\mathbf{S})$, and $d_{\text{head}}$ is the scaling factor determined by the attention head dimension. Different from traditional linear, fixed operators, these learned attention weights naturally serve as flexible and data-driven inverse operators, adaptively integrating spatiotemporal sensor properties and temporal signal dynamics to estimate latent variables.
@@ Appendix C Related Work on Brain Foundation Models
Appendix D Pretraining Dataset Description
@@
Appendix E Downstream Finetuning Dataset Description
@@
  • PhysioNet-MI[56]: The PhysioNet-MI dataset includes EEG recordings ( 64 channels, 160 Hz ) from 109 subjects performing 4 motor imagery and movement tasks. A total of 19571 4-second samples were used for classification to predict the type of task.
@@
  • AD65[41]: The AD65 dataset includes EEG recordings ( 19 channels, 500 Hz ) from 36 subjects with Alzheimer’s disease and 29 healthy subjects during resting state. A total of 7013 10-second samples were used for classification to predict the presence of Alzheimer’s disease.
Appendix F Baseline Models Description
@@
Appendix G Implementation Details
We trained BrainTokenizer using 2-second segments. For training BrainOmni, we inputted 30-second data segments to allow the model to capture longer temporal dependencies. During the segmented tokenization process, we set the overlap ratio between windows to 25% to incorporate partial contextual information. The training was conducted on 16 A100 GPUs, using the AdamW optimizer and a warmup-cosine-decay learning rate scheduler with a warmup proportion of 10%. BrainTokenizer was trained for 16 epochs with a total batch size of 512 per update step and a maximum learning rate of 2e-4. BrainOmni was trained for 32 epochs with a total batch size of 256 per update step and a maximum learning rate of 4e-4.
Appendix H Ablation Study
H.1 Latent Source Activity Estimation
As mentioned in the Introduction, modeling at the electrode level can encounter issues such as low information density and redundant information in adjacent channels. The latent source activity modeling process involves compressing the sequence length, which, while condensing semantic features, may also introduce some loss of detail. To further demonstrate the advantages of modeling source activity compared to modeling at the electrode level, we trained a model that integrates spatial features through self-attention between electrodes, which would not involve loss of information. To be more specific, the BrainTokenizer of this model replaces the cross attention layer with self-attention layer. Channel random masking and additive noise are also applied before the self-attention layer in the encoder part of the feature input. In the decoder part, the masked channels are replaced with a mask token, and the waveform of the masked part is recovered by a self-attention layer.
Table 7: Balanced accuracy results of ablation study for latent source activity estimation (LSAE). For the model with LSAE, the BrainTokenizer compresses the channels into 16 channels uniformly.
TUAB PD31 AD65 ASD74 SomatoMotor
tiny w/o NAE 0.814 $\pm$ 0.002 0.595 $\pm$ 0.051 0.792 $\pm$ 0.034 0.605 $\pm$ 0.052 0.805 $\pm$ 0.127
tiny w/ NAE 0.816 $\pm$ 0.001 0.658 $\pm$ 0.062 0.800 $\pm$ 0.035 0.606 $\pm$ 0.019 0.787 $\pm$ 0.112
From the table above, we can see that even though models using latent source activity estimation may lose some information during the sequence length compression process, their performance on EEG datasets such as TUAB (23), PD31 (32), ad65 (19), asd74 (306) and SomatoMotor (372) still exceeds that of models built at the electrode level. The numbers in parentheses represent the number of channels in each dataset. For the SomatoMotor (372) dataset, even with a compression ratio of around 23 times, the performance of the model using LSAE is only 0.018 lower than that of the uncompressed model. Moreover, since the computational complexity of the Transformer model grows quadratically with sequence length, models using LSAE have significantly higher computational efficiency compared to those that do not use it. This demonstrates that the process of LSAE can eliminate redundant information between channels while retaining core semantic features. It also illustrates the benefits of projecting all data from the electrode level into a unified feature space for modeling.
H.2 Freeze Backbone for Downstream
In Table 8, we report the BACC metrics of BrainOmni and the baseline models under both full fine-tuning and frozen settings on TUAB, TUEV, and AD65 datasets. In TUAB and AD65, the performance of BrainOmni under the weight freezing scenario is comparable to or even higher than the metrics of the baseline models with full fine-tuning. This indicates that BrainOmni has learned the pattern features of EMEG signals with better generalization capability through unsupervised training.
Table 8: Results for freeze the backbone model when downstream training. Results with full finetuning also listed for comparison. BACC reported.
TUAB TUEV AD65
LaBraM Freeze 0.788 $\pm$ 0.005 0.394 $\pm$ 0.030 0.699 $\pm$ 0.025
Full Finetune 0.807 $\pm$ 0.003 0.596 $\pm$ 0.016 0.706 $\pm$ 0.012
CBraMod Freeze 0.768 $\pm$ 0.001 0.398 $\pm$ 0.067 0.633 $\pm$ 0.003
Full Finetune 0.773 $\pm$ 0.001 0.416 $\pm$ 0.042 0.649 $\pm$ 0.006
BrainOmni_tiny Freeze 0.806 $\pm$ 0.001 0.464 $\pm$ 0.021 0.772 $\pm$ 0.046
Full Finetune 0.816 $\pm$ 0.001 0.597 $\pm$ 0.019 0.800 $\pm$ 0.035
BrainOmni_base Freeze 0.811 $\pm$ 0.001 0.489 $\pm$ 0.016 0.780 $\pm$ 0.058
Full Finetune 0.813 $\pm$ 0.007 0.618 $\pm$ 0.034 0.806 $\pm$ 0.053
Appendix I Hyperparameters for model
@@
Quantize optimize method "ema"
Latent source numberz 16
@@
Total batch per update 64
@@
Appendix J Attention Weights of Cross Attention in the BrainTokenizer
Recall the content in Appendix B.2. The cross-attention mechanism in BrainTokenizer is inspired by the inverse operator in the Backward Solution. The attention weights can actually be regarded as a parameterised linear inverse matrix, representing the degree of influence of each channel on each latent source activity. From Fig. 6, we can see that each source variable in our model has automatically learned a hierarchical structure. It is evident that each source variable focuses on extracting features from specific scalp regions. Moreover, in the multi-head attention mechanism, the attention of the first head is relatively concentrated, while the subsequent heads progressively expand the range of regions.
Figure 6: Visualization of the topographic maps of cross-attention for each channel in BrainTokenizer. The model’s 16 queries are displayed in 8 columns on the top and bottom, with the first eight and the last eight queries shown separately. The 4 rows represent the attention weights of each head in the multi-head cross-attention mechanism.

### v3

@@ 5.1 Effectiveness of Sensor Encoder
An ablation study is conducted to assess the effectiveness of the proposed Sensor Encoder in modelling the spatial characteristics of sensors. Apart from directly removing the sensor embedding, we also develop a pure temporal version of BrainTokenizer where multi-channel EMEG data is treated as uncorrelated single-channel data. As shown in Table 6, the standard BrainOmni consistently outperforms the pure temporal model. And the exclusion of sensor embedding significantly undermines the downstream performance, especially on challenging MEG and EMEG datasets, which demonstrates the effectiveness of the proposed Sensor Encoder.
5.2 Impact of Loss Items in BrainTokenizer
The BrainTokenizer training loss contains four components: time-domain loss, frequency-domain loss, Pearson correlation loss, and RVQ commitment loss. Given that the time-domain and commitment losses are essential for training, we conducted an ablation study for how frequency loss and PCC loss influence the reconstruction performance. Results are shown in Table 7. With only time-domain and commitment loss, the decoder outputs collapsed into near‑constant signals and got poor performance. Introducing the frequency‑domain loss greatly improved spectral detail (lower MAE/AMP/PHASE), but led to spurious high‑frequency impulses. The PCC term mitigates these impulses by aligning overall waveform trends, yet on its own provides weaker spectral reconstruction. All four components were included in the BrainTokenizer training, suppressing the impulse while preserving as much frequency-domain detail as possible.
Table 7: Reconstruction results of BrainTokenizer w or w/o freq loss and PCC loss. Total loss denotes the sum of all four loss items.
Freq loss PCC loss MAE $\downarrow$ AMP $\downarrow$ PHASE $\downarrow$ PCC $\uparrow$ Commitment $\downarrow$ Total Loss $\downarrow$
0.799 0.256 2.154 0.002 5.01e-7 4.207
0.471 0.160 1.583 0.716 1.64e-4 2.517
0.462 0.167 1.863 0.751 1.83e-4 2.703
0.480 0.165 1.495 0.723 1.48e-4 2.626
(a) Latent source variables
(b) Learning curve
(c) Different codebook layers
Figure 3: (a) Trend of L1 loss over the number of latent source variables. (b) The training loss and validation loss curves during the training phase of BrainTokenizer. (c) The accuracy curves for parallel mask prediction on the labels of each codebook layer during the training phase of BrainOmni on the training set.
5.3 Number of Latent Source Variables
To investigate the effect of the number of latent source activities, we trained variants of BrainTokenizer with different numbers of latent source variables. Results are shown in Fig. 3(a). As the number of source activities increases, the EMEG signal reconstruction loss gradually decreases. When the number of source variables reaches beyond 16, the trend of loss reduction weakened with further increases in the number of latent source activities. To balance the amount of information retained and computational efficiency, we chose 16 as the number of latent source variables.
@@ 5.4 Pretraining Stability and Codebook Contribution
To analyse the training stability when incorporating diverse devices and signals into pretraining, we plot the training curves in Fig. 3(b). Overall, BrainTokenizer converges relatively quickly, and the validation loss remains consistent with the training loss. This demonstrates that the proposed BrainTokenizer training framework handle different devices and signal types well. Fig. 3(c) plots the mask prediction performance for each RVQ layer. It can be seen that the first codebook aggregates richer semantic information and achieves higher accuracy in mask prediction, while the prediction accuracy decreases progressively for the subsequent lower-level codebooks.
@@ 5.5 Visualisation of Reconstruction Results
Fig. 4 illustrates the reconstruction performance of BrainTokenizer. We compare the waveforms and topographic maps before and after reconstruction to show the model’s ability to capture the spatiotemporal electromagnetic fields of EEG and MEG in both temporal and spatial domains. The results show that the model effectively preserves the main trends and finer details of the waveforms, while smoothing out high-frequency noise spikes. In terms of spatial representation, the reconstructed topographic maps maintain the original activation patterns and structural integrity. Notably, the model also achieves strong reconstruction performance on previously unseen devices and samples, demonstrating robust generalisation across different recording setups.
@@ 6 Conclusion
In this paper, we introduce BrainOmni, the first brain foundation model that generalises across heterogeneous EEG and MEG signals. To model signals from different devices and modalities, we propose BrainTokenizer, which infers spatiotemporal patterns of brain activity from the observed EMEG signals and generates quantized discrete tokens. To address the inherent heterogeneity arising from various recording devices, we develop a flexible Sensor Encoder that leverages physical sensor properties. Through large-scale self-supervised joint pretraining on EEG and MEG data, BrainOmni significantly exceeds the performance of existing foundation models and state-of-the-art task-specific baselines across various downstream tasks, and demonstrates effective generalisation capability to unseen devices and modalities. Furthermore, extensive ablation analysis highlights the consistent benefits of joint EMEG pretraining, underscoring the advantage of unified modelling strategies. We believe BrainOmni represents an important step towards building versatile and scalable foundation models for neural recordings, opening new pathways for unified brain signal representation learning.
7 Acknowledgements
This work was supported by the New Generation Artificial Intelligence – National Science and Technology Major Project of China (Grant No. 2025ZD0121803) and the National Natural Science Foundation of China (Grant No. 62476151).
@@ Appendix A Broader Impacts and Limitations
As the first foundation model enabling unified analysis across EEG and MEG modalities, BrainOmni represents a crucial step toward generalizable and scalable neural representation learning, offering substantial application potential in neuroscience research, health monitoring, and clinical diagnosis. Additionally, the model’s enhanced generalisation on heterogeneous devices could reduce the performance degradation typically associated with device and dataset variability. While it may not always outperform domain-specific models tailored to one dataset or one task, it offers broader coverage and generalisability across datasets, tasks, and modalities – a key advantage for scalable and reusable neuro-AI systems.
@@ B.1 Source Current Estimation
$V(\mathbf{r}_{m})=\int_{\Omega}G(\mathbf{r}_{m},\mathbf{r}^{\prime})\nabla\cdot\mathbf{J}_{p}(\mathbf{r}^{\prime})\,d\mathbf{r}^{\prime}$ (9)
@@
$\mathbf{B}(\mathbf{r}_{m})=\frac{\mu_{0}}{4\pi}\int_{\Omega}\frac{\mathbf{J}_{p}(\mathbf{r}^{\prime})\times(\mathbf{r}_{m}-\mathbf{r}^{\prime})}{|\mathbf{r}_{m}-\mathbf{r}^{\prime}|^{3}}\,d\mathbf{r}^{\prime}$ (10)
@@
The inverse problem – estimating the sources $\mathbf{j}(t)$ from the measurements $\mathbf{y}(t)$ – is ill-posed and requires regularization. Minimum-norm estimation solves this by seeking the source configuration with the smallest $\ell_{2}$-norm that still explains the data:
@@ B.2 Model Explanation From the Perspective of Neuroscience
Inspired by this conceptual framework, our proposed BrainTokenizer closely mirrors these two steps. Specifically, the BrainTokenizer (excluding the RVQ quantisation step) corresponds directly to the backward solution stage, where it maps the raw EMEG signals and sensor device layout into latent source variables through a parametric module. Recall the classical backward solution,
@@
which relies on a fixed linear inverse operator $\mathbf{W}$ constructed from a leadfield matrix $\mathbf{L}$ and a noise covariance matrix $\mathbf{C}_{n}$. While in our architecture, we use the attention weights in cross-attention mechanisms to construct a time-varying parametric inverse operator $\mathbf{W}_{\mathbf{\theta}}(\mathbf{Z}_{\text{time}},\mathbf{L},\mathbf{S})$, where the operator is related to sensor prosperity and the measured data since the different fields recorded outside may indicate that there is a change in the pattern of neuro activity, formulated as:
@@
where $\mathbf{Q}$ is a set of learnable query embeddings that dynamically aggregate temporal and spatial information, $\mathbf{K}$ is the combination of temporal representation $\mathbf{Z}_{\text{time}}$ from SEANet encoder and sensor embeddings $\mathbf{V}(\mathbf{L},\mathbf{S})$, and $d_{\text{head}}$ is the scaling factor determined by the attention head dimension. Different from traditional linear, fixed operators, these learned attention weights naturally serve as flexible and data-driven inverse operators, adaptively integrating spatiotemporal sensor properties and temporal signal dynamics to estimate latent variables. Appendix G provides a visualisation for attention weights of cross attention.
@@ Appendix C Related Work on Brain Foundation Models
Appendix D Implementation Details
We trained BrainTokenizer using 2-second segments. For training BrainOmni, we inputted 30-second data segments to allow the model to capture longer temporal dependencies. During the segmented tokenization process, we set the overlap ratio between windows to 25% to incorporate partial contextual information. The training was conducted on 16 A100 GPUs, using the AdamW optimizer and a warmup-cosine-decay learning rate scheduler with a warmup proportion of 10%. BrainTokenizer was trained for 16 epochs with a total batch size of 512 per update step and a maximum learning rate of 2e-4. BrainOmni was trained for 32 epochs with a total batch size of 256 per update step and a maximum learning rate of 4e-4. The BrainTokenizer training took approximately 11 hours, and BrainOmni required about 14 hours for the tiny model and 18 hours for the base model.
For downstream evaluation, all models follow a unified training pipeline. The output embeddings from each model are first average pooled along the temporal dimension, then flattened across remaining dimensions to serve as feature, which are subsequently fed into a two-layer MLP for classification. Taking BrainOmni as an example, for a batch of samples, the model outputs a embedding with shape $(B,C^{\prime},T,D)$, where $B$ represents for batch size, $C^{\prime}$ the number of compressed channels, $T$ temporal sequence length, and $D$ embedding dimension. The embeddings is firstly average pooled over $T$, and then flattened to $(B,C^{\prime}*D)$, before fed into the MLP classifier.
The evaluation was conducted in a cross-subject five-fold cross-validation paradigm. For datasets without an official split (i.e., datasets except TUAB and TUEV), all subjects were divided into five subsets, and five experiments were conducted in rotation. In each experiment, three folds were used for training, one for validation, and one for testing, ensuring that every subset served as the test set once, thereby providing a more comprehensive assessment of the model’s performance. For TUAB and TUEV which provided an official train/eval split, we kept the eval set for all testing, and split the official training subjects equally into five folds, and in each runs use four folds for training and remaining fold for validation. All experiments were run under two random seeds (42 and 3407) and seed 42 is used for data splitting. Experiments were conducted with learning rates of $3\times 10^{-6},1\times 10^{-5},3\times 10^{-5}]$ for each model, and the learning rate yielding the highest result was selected to obtain the best performance of each model. All other hyperparameters remained consistent across experiments (specific values are provided in Table 12). On a single A100 GPU, the training throughput was approximately 60 samples/sec, while inference achieved roughly 90 samples/sec.
Appendix E Preprocess Details
The power-spetral density-based bad-channel detection is implemented as Alg. 1.
Input : raw_data, $\textit{threshold}=10$
Output : bad_channels
1. Compute per-channel PSD over the full recording:
$\mathrm{PSD}\leftarrow\textsc{ComputePSD}(\textit{raw\_data}).\textit{data}$
2. Stabilise and log-transform:
$L\leftarrow\log(\mathrm{PSD}+10^{-16})$
3. Compute pairwise distances between channel spectra:
for $i\leftarrow 1$ to $C$ do
   for $j\leftarrow 1$ to $C$ do
     $\mathrm{dist}[i,j]\leftarrow L[i]-L[j]$
     end for
     end for
    4. Mean distance per channel:
     $m[i]\leftarrow\mathrm{mean}_{j}\big(\mathrm{dist}[i,j]\big)$ for $i=1,\dots,C$
    5. Identify outliers via IQR:
     $Q_{1}\leftarrow\mathrm{percentile}_{25}(m)$,  $Q_{3}\leftarrow\mathrm{percentile}_{75}(m)$
     $\mathrm{IQR}\leftarrow Q_{3}-Q_{1}$
     $\mathrm{upper}\leftarrow Q_{3}+\textit{threshold}\cdot\mathrm{IQR}$,  $\mathrm{lower}\leftarrow$ $Q_{1}-\textit{threshold}\cdot\mathrm{IQR}$
     $\textit{bad\_channels}\leftarrow\{\,\mathrm{ch}_{i}\mid m[i]>\mathrm{upper}\ \vee\ m[i]<\mathrm{lower}\,\}$
Algorithm 1 Power-spectral-density-based bad-channel detection
To address the reference inconsistency caused by different recording setups, we first compute a global average reference by taking the mean across all channels, and then subtract this channel-average signal from each channel waveform, effectively aligning all channels to the same virtual reference. This protocol is applied regardless of whether a reference channel is unavailable or the signals have already been referenced, and it is performed both at the recording level and separately for each signal type.
Appendix F Ablation Study
F.1 Latent Source Activity Estimation
As mentioned in the Introduction, modelling at the electrode level can encounter issues such as low information density and redundant information in adjacent channels. The latent source activity modelling process involves compressing the sequence length, which, while condensing semantic features, may also introduce some loss of detail. To further demonstrate the advantages of modelling source activity compared to modelling at the electrode level, we trained a model that integrates spatial features through self-attention between electrodes, which would not involve loss of information. To be more specific, the BrainTokenizer of this model replaces the cross attention layer with self-attention layer. Channel random masking and additive noise are also applied before the self-attention layer in the encoder part of the feature input. In the decoder part, the masked channels are replaced with a mask token, and the waveform of the masked part is recovered by a self-attention layer. Results are show in Table 8.
Table 8: Balanced accuracy results of ablation study for latent source activity estimation (LSAE)
TUAB PD31 AD65 ASD74 SomatoMotor
tiny w/o NAE 0.813 $\pm$ 0.003 0.627 $\pm$ 0.158 0.768 $\pm$ 0.056 0.606 $\pm$ 0.088 0.801 $\pm$ 0.078
tiny w/ NAE 0.819 $\pm$ 0.004 0.736 $\pm$ 0.116 0.795 $\pm$ 0.030 0.621 $\pm$ 0.048 0.863 $\pm$ 0.128
We can see that even though models using latent source activity estimation may lose some information during the sequence length compression process, their performance on datasets such as TUAB (23), PD31 (32), AD65 (19), ASD74 (306) and SomatoMotor (372) still exceeds that of models built at the electrode level. The numbers in parentheses represent the number of channels in each dataset. Moreover, since the computational complexity of the Transformer model grows quadratically with sequence length, models using LSAE have significantly higher computational efficiency compared to those that do not use it. This demonstrates that the process of LSAE can eliminate redundant information between channels while retaining core semantic features. It also illustrates the benefits of projecting all data from the electrode level into a unified feature space for modelling.
F.2 Freeze Backbone for Downstream
In Table 9, we report the BACC metrics of BrainOmni and the baseline models under both full fine-tuning and frozen settings on TUAB, TUEV, and AD65 datasets. In TUAB and AD65, the performance of BrainOmni under the weight freezing scenario is comparable to or even higher than the metrics of the baseline models with full fine-tuning. This indicates that BrainOmni has learned the pattern features of EMEG signals with better generalisation capability through unsupervised training.
Table 9: Downstream results with frozen backbone vs. full finetuning (BACC).
TUAB TUEV AD65
LaBraM Freeze 0.800 $\pm$ 0.004 0.432 $\pm$ 0.024 0.722 $\pm$ 0.039
Full Finetune 0.816 $\pm$ 0.006 0.588 $\pm$ 0.017 0.711 $\pm$ 0.060
CBraMod Freeze 0.774 $\pm$ 0.001 0.345 $\pm$ 0.014 0.620 $\pm$ 0.057
Full Finetune 0.808 $\pm$ 0.007 0.525 $\pm$ 0.021 0.681 $\pm$ 0.040
BrainOmni_tiny Freeze 0.800 $\pm$ 0.002 0.460 $\pm$ 0.021 0.774 $\pm$ 0.048
Full Finetune 0.819 $\pm$ 0.004 0.603 $\pm$ 0.024 0.795 $\pm$ 0.030
BrainOmni_base Freeze 0.809 $\pm$ 0.003 0.480 $\pm$ 0.015 0.771 $\pm$ 0.062
Full Finetune 0.819 $\pm$ 0.005 0.622 $\pm$ 0.028 0.828 $\pm$ 0.030
Appendix G Attention Weights of Cross Attention in the BrainTokenizer
Recall the content in Appendix B.2, the cross-attention mechanism in BrainTokenizer is inspired by the inverse operator in the Backward Solution. The attention weights can actually be regarded as a parameterised linear inverse matrix, representing the degree of influence of each channel on each latent source activity. From Fig. 6, we can see that each source variable in our model has automatically learned a hierarchical structure. It is evident that each source variable focuses on extracting features from specific scalp regions. Moreover, in the multi-head attention mechanism, the attention of the first head is relatively concentrated, while the subsequent heads progressively expand the range of regions.
Figure 6: Visualisation of the topographic maps of cross-attention for each channel in BrainTokenizer. The model’s 16 queries are displayed in 8 columns on the top and bottom, with the first eight and the last eight queries shown separately. The 4 rows represent the attention weights of each head in the multi-head cross-attention mechanism.
Appendix H Pretraining Dataset Description
@@
Appendix I Downstream Finetuning Dataset Description
@@
  • PhysioNet-MI[58]: The PhysioNet-MI dataset includes EEG recordings (64 channels, 160 Hz) from 109 subjects performing 4 motor imagery and movement tasks. A total of 9837 4-second samples were used for classification to predict the type of task.
@@
  • AD65[43]: The AD65 dataset includes EEG recordings (19 channels, 500 Hz) from 36 subjects with Alzheimer’s disease and 29 healthy subjects during resting state. A total of 5349 10-second samples were used for classification to predict the presence of Alzheimer’s disease.
  • MEG-MMI[40]: The MEG-MMI dataset includes MEG recordings (269 magnetometer channels, 1200 Hz) from 29 adolescents with major depression and 22 healthy subjects during mood induction tasks. A total of 1770 30-second samples were used for classification to predict the presence of major depression.
Appendix J Baseline Models Description
@@
  • FAMED[31]: FAMED is a non-pretrained model for MEG-based epilepsy analysis, built on SCSE-ResNet with dilated convolutions.
@@
Appendix K Hyperparameters for model
@@
Latent source number 16
@@
Total batch per update 128